OpenStudy (anonymous):

Can someone please check my work? - ABCD is equal to QRST. - The measure of angle A is x-10 & the measure of angle Q is 2x-30 -- What is the Measure of angle A?? I got 20 becuase I think that A and Q would be equal. and 20-x is 10 as is 2x-30.

OpenStudy (anonymous):

Assuming A = Q, you can set the equations equal to each other. x-10 = 2x - 30 Subtract x from each side and add 30 to each side. 20 = x Now, to get the angle of A, you have to plug 20 into the equation for A. A = x-10 A = 20-10 A = 10

OpenStudy (anonymous):

ohhh. Ok. I gotcha! I was on the rightish track!! lol Thanks!! :D